# Emergence of the Primacy Effect in Structured State-Space Models

### v1

@@ 6.3 Implications for Computational Neuroscience
Finally, the present study does not offer an explanation for various well-documented properties of the psychological primacy effect. For instance, the retention of initial study items is thought to be achieved, or at least facilitated, by rehearsal—repetitive mental or verbal recitation of observed items before requested recall/verification (Rundus,, 1971; Marshall and Werder,, 1972; Glenberg et al.,, 1980)—for which the models studied here lack an analogous mechanism. In addition, the models operated on a unitless discrete time scale, to which the continuous-time representation of the SSM was aligned. As a result, they are unsuitable for investigating physical-time effects, such as the impact of retention intervals between study-item presentation and verification (Cornell and Bergstrom,, 1983; Wright et al.,, 1985; Neath,, 1993) or variations in inter-item intervals (Neath and Knoedler,, 1994). Future studies may address this latter issue by employing more biologically realistic neural network models; as noted above, the SSM is closely related to spiking neural networks, which are capable of simulating physical-time dynamics and incorporating biologically inspired connectivity structures (Voelker and Eliasmith,, 2018).

### v2

@@ 5.3 Implications for Computational Neuroscience
Finally, the present study does not offer an explanation for various well-documented characteristics of the psychological primacy effect. For instance, the retention of initial study items is thought to be achieved, or at least facilitated, by rehearsal—repetitive mental or verbal recitation of observed items before requested recall/verification (Rundus,, 1971; Marshall and Werder,, 1972; Glenberg et al.,, 1980)—for which the models studied here lack an analogous mechanism.
In addition, the models operated on a unitless discrete time scale, to which the continuous-time representation of the SSM was aligned. As a result, they are unsuitable for investigating physical-time effects, such as the impact of retention intervals between study-item presentation and verification (Cornell and Bergstrom,, 1983; Wright et al.,, 1985; Neath,, 1993) or variations in inter-item intervals (Neath and Knoedler,, 1994). This limitation could be addressed by exploring spiking neural networks equivalent the SSM. As noted above, the SSM can be converted into a spiking neural network, which in turn enables the simulation of the physical-time effects as well as the incorporation of biologically inspired connectivity structures (Voelker and Eliasmith,, 2018).